OpenStudy (anonymous):

f(x) = x + 8 and g(x) = x – 4. What is f(x) • g(x)?

OpenStudy (inkyvoyd):

is that a multiplecation dot, or a composite function dot?

OpenStudy (anonymous):

multiplication

OpenStudy (inkyvoyd):

Then, it's (x+8)(x-4)

OpenStudy (anonymous):

a.)x2 – 4x – 32 b.)x2 + 4x – 32 c.)x2 – 12x + 4 d.)x2 + 4x + 4 it has to be one of these thats why im lost lol

OpenStudy (anonymous):

f(x) = x + 8 and g(x) = x – 4. f(x) • g(x)= (x + 8)(x – 4)..simplify more this...

OpenStudy (anonymous):

\[\Large (x+8)(x-4)=x^2-4x+8x-32=x^2+4x-32\]

OpenStudy (anonymous):

xx+8x-4x-32=x^2 +4x-32